# Event-Based Simulation of Stochastic Memristive Devices for Neuromorphic Computing

### v1

@@ V Conclusions and future work
We have introduced an event-based modelling strategy for memristive devices by extending previous generic modelling approaches for memristors. We introduced an event-based modelling strategy based on Poisson processes which addressed errors associated with discrete time approximation in existing models, as well as being natively suited to neuromorphic applications. We proposed a volatility state variable, to allow for the capture of memory-dependent switching behaviour, motivating this from a physical perspective for various memristive device types. We have also presented a resistive drift dataset, which we use to obtain parameters for an example of the model applied to a real device, along with a new linear conductance model of drift in the devices. We demonstrated the ability of the event-based model to replicate the behaviour under zero bias. Future work may include extending the modelling approaches described to additional real data from different memristive device types, under a variety of additional conditions, as well as exploration of the applications of the model for event-based neuromorphic systems and simulations, especially spiking neural networks implemented in hardware.

### v2

@@ V Conclusion
We have introduced an event-based modelling strategy for memristive devices where time is the main computational resource. In such systems, event-based simulation is essential, and use of a volatility state variable can help to make clear routes to controlling device behaviour through the spacing of input to the memristive device over time, leading to the possibility of the emulation of different forms of neuronal dynamics, dependent on the stability of the memristive device in question. Our framework, based on multiple parallel Poisson processes, addresses errors associated with discrete time approximation in existing models, as well as being natively suited to event-driven neuromorphic applications. We proposed a volatility state variable, to allow for the capture of transient switching dynamics, motivating this from a physical perspective for various memristive device types. We have also presented a resistive drift dataset for resistive drift in titanium dioxide memristors, which we used to obtain parameters for an example of the model applied to a real device, along with proposing a new linear conductance model of drift in the devices, motivated by a model of the connection between the internal state and the observed resistance. We demonstrated the ability of the event-based model to replicate the resistive drift behaviour under zero bias following the development of the Poisson rate parameter model and volatility model for the titanium dioxide memristors, and subsequent parameter fitting. Finally, we examined the modelling framework from the perspective of neuromorphic computing, studying possible requirements for parameters of future devices that may enable frequency-dependent switching, and replication of action potentials using memristive devices. We introduced the notions of stable and unstable memristors, based on their switching rate in low bias conditions, and demonstrated how our proposed modelling approach based on the volatility state variable allows for better understanding and control of the neuromorphic behaviours. Future work can include applying the modelling framework described to additional device types, specifying probabilistic switching models, volatility models, and readout equations.